In a stairway there are 36 steps of equal height. If each step were 1 inch higher,
there would have been 32steps. How high is each step?

Answer by Theo(3460)

36 steps of size x
32 steps of size (x+1)

since the vertical distance is the same, this means that:

36*x = 32*(x+1)

remove parentheses to get:

36*x = 32*x + 32

subtract 32*x from both sides to get:

4*x = 32

divide both sides fo 4 to get:

x = 8

x + 1 = 9

36*8 = 288 feet.

32*9 = 288 feet.

same elevation so we should be good.